### v1

@@ Abstract
In this paper, we introduce a multi-robot system that integrates mapping, localization, and task and motion planning (TAMP) enabled by 3D scene graphs to execute complex instructions expressed in natural language. Our system builds a shared 3D scene graph incorporating an open-set object-based map, which is leveraged for multi-robot 3D scene graph fusion. This representation supports real-time, view-invariant relocalization (via the object-based map) and planning (via the 3D scene graph), allowing a team of robots to reason about their surroundings and execute complex tasks. Additionally, we introduce a planning approach that translates operator intent into Planning Domain Definition Language (PDDL) goals using a Large Language Model (LLM) by leveraging context from the shared 3D scene graph and robot capabilities. We provide an experimental assessment of the performance of our system on real-world tasks in large-scale, outdoor environments.
@@ 3 Multi-Robot 3D Scene Graph Fusion, Position Estimation, and Relocalization
Single-robot Mapping. 3D scene graphs provide a rich, hierarchical structure which is useful for downstream planning and can be built in real-time onboard a robot. Hydra’s scene graphs contain a detailed mesh of the environment, an object layer, a places layer representing 3D free space, and a region layer representing larger areas of consistent semantic meaning (e.g., road or sidewalk). We add a layer of surface places, proposed in [Ray24isrr-tamp], to represent navigable 2D space. The surface places each represent a section of the mesh with consistent semantic labels (e.g., grass or rocks), which are computed by partitioning the mesh vertices according to their distance from a set of generating place centers. This process repeats incrementally in the Hydra backend, mimicking the behavior of Lloyd’s algorithm [lloyds] for computing a Voronoi partition and resulting in evenly spaced convex polygons. The surface places layer has edges connecting neighboring places, so a path in the subgraph of traversable surface places corresponds to a reference path for a ground robot. We use ROMAN [peterson2025roman] to build an open-set object map to detect loop closures and enable relocalization to a prior map. ROMAN uses FastSAM [peterson2025roman] to segment color images and then uses depth measurements to track these segments as 3D objects. Additionally, CLIP embeddings [radford2021clip] are computed to give objects an open-set semantic descriptor. ROMAN formulates a robust optimization problem for object matching using geometry, shape, and semantic similarity to associate objects between maps. The key insight is that object-based maps are more view-invariant than image features, enabling loop closures across extreme viewpoint changes.
@@ 4 Language-Grounded Planning
Figure 2: Example of LLM prompt and response including snippets for the descriptions of the task, domain, robot, examples, and instructions.
@@ 5.2 Ablations
Scene Graph Fusion. To understand the performance of the scene graph fusion beyond the qualitative results from the end-to-end experiments, we examined the accuracy of our multi-robot scene graph fusion on the Kimera-Multi dataset [Tian23iros-kimeraMulti].We do not complete the ablations during the real-time experiments at Camp Buckner due to the lack of ground truth. Specifically, we used the hybrid and outdoor sequences, and for each sequence, we used the data for the acl_jackal2, sparkal1, and sparkal2 robots. To showcase the benefits of ROMAN, we compared against a competitive visual loop-closure baseline that uses SALAD [izquierdo2024salad] for place descriptors, SuperPoint [detone2018superpoint] for local features, and LightGlue [lindenberger2023lightglue] for feature matching. We report the Root Mean Square Error (RMSE) of the Absolute Trajectory Error (ATE) of the optimized robot trajectories, along with the precision, recall, and Intersection over Union (IoU) of the objects in the fused scene graph as defined in [Chang23iros-HydraMulti] with a tolerance of $5m$. The results are shown in Table 2. These metrics validate the improved accuracy of ROMAN when integrated with Hydra-Multi as compared to traditional visual place recognition based on visual place descriptors and feature matching. While loop closures from ROMAN resulted in a slightly higher ATE RMSE, the optimized scene graphs tended to have higher global consistency, resulting in improvements in object IoU, precision, and recall. Additionally, ROMAN requires minimal data storage and transmission for computing loop closures. We observed better performance when ROMAN has access to gravity-aligned odometry (e.g., from IMU), as demonstrated by the ATE RMSE result of $6.91m$ reported in [peterson2025roman] on the same $hybrid$ data sequences. Furthermore, by experimental design, much of the trajectory overlap occured when robots were facing similar directions where visual feature matching methods excel.
@@ 7 Conclusions
While the results are promising, this work opens several directions for future research. For example, the robot is currently constrained to operating within the previously fused map, which limits the system from handling exploration tasks or adapting to changes in the environment. Further, the language grounding does not yet exploit the open-set nature of the object-based map to support more open-ended tasks. These are exciting directions for future research, with the potential to enable robots to complete a wider variety of tasks.

### v2

@@ Abstract
In this paper, we introduce a multi-robot system that integrates mapping, localization, and task and motion planning (TAMP) enabled by 3D scene graphs to execute complex instructions expressed in natural language. Our system builds a shared 3D scene graph incorporating an open-set object-based map, which is leveraged for multi-robot 3D scene graph fusion. This representation supports real-time, view-invariant relocalization (via the object-based map) and planning (via the 3D scene graph), allowing a team of robots to reason about their surroundings and execute complex tasks. Additionally, we introduce a planning approach that translates operator intent into Planning Domain Definition Language (PDDL) goals using a Large Language Model (LLM) by leveraging context from the shared 3D scene graph and robot capabilities. We provide an experimental assessment of the performance of our system on real-world tasks in large-scale, outdoor environments. A supplementary video is available at https://youtu.be/8xbGGOLfLAY.
@@ 3 Multi-Robot 3D Scene Graph Fusion, Position Estimation, and Relocalization
Single-robot Mapping. 3D scene graphs provide a rich, hierarchical structure which is useful for downstream planning and can be built in real-time onboard a robot. Hydra’s scene graphs contain a detailed mesh of the environment, an object layer, a places layer representing 3D free space, and a region layer representing larger areas of consistent semantic meaning (e.g., road or sidewalk). We add a layer of surface places, proposed in [Ray24isrr-tamp], to represent navigable 2D space. The surface places each represent a section of the mesh with consistent semantic labels (e.g., grass or rocks), which are computed by partitioning the mesh vertices according to their distance from a set of generating place centers. This process repeats incrementally in the Hydra backend, mimicking the behavior of Lloyd’s algorithm [lloyds] for computing a Voronoi partition and resulting in evenly spaced convex polygons. The surface places layer has edges connecting neighboring places, and the resulting graph of places can be used for path planning. We use ROMAN [peterson2025roman] to build an open-set object map to detect loop closures and enable relocalization to a prior map. ROMAN uses FastSAM [peterson2025roman] to segment color images and then uses depth measurements to track these segments as 3D objects. Additionally, CLIP embeddings [radford2021clip] are computed to give objects an open-set semantic descriptor. ROMAN formulates a robust optimization problem for object matching using geometry, shape, and semantic similarity to associate objects between maps. The key insight is that object-based maps are more view-invariant than image features, enabling loop closures across extreme viewpoint changes.
@@ 4 Language-Grounded Planning
Figure 2: Example of LLM prompt and response including snippets for the descriptions of the translation task, PDDL domain, robot capabilities, in-context examples, and instructions for the robots.
@@ 5.2 Ablations
Scene Graph Fusion. To understand the performance of the scene graph fusion beyond the qualitative results from the end-to-end experiments, we examined the accuracy of our multi-robot scene graph fusion on the Kimera-Multi dataset [Tian23iros-kimeraMulti] [^1] We do not complete the ablations during the real-time experiments at Camp Buckner due to the lack of ground truth. Specifically, we used the hybrid and outdoor sequences, and for each sequence, we used the data for the acl_jackal2, sparkal1, and sparkal2 robots. To showcase the benefits of ROMAN, we compared against a competitive visual loop-closure baseline that uses SALAD [izquierdo2024salad] for place descriptors, SuperPoint [detone2018superpoint] for local features, and LightGlue [lindenberger2023lightglue] for feature matching. We report the Root Mean Square Error (RMSE) of the Absolute Trajectory Error (ATE) of the optimized robot trajectories, along with the precision, recall, and Intersection over Union (IoU) of the objects in the fused scene graph as defined in [Chang23iros-HydraMulti] with a tolerance of $5m$. The results are shown in Table 2. These metrics validate the improved accuracy of ROMAN when integrated with Hydra-Multi as compared to traditional visual place recognition based on visual place descriptors and feature matching. While loop closures from ROMAN resulted in a slightly higher ATE RMSE, the optimized scene graphs tended to have higher global consistency, resulting in improvements in object IoU, precision, and recall. Additionally, ROMAN requires minimal data storage and transmission for computing loop closures. We observed better performance when ROMAN has access to gravity-aligned odometry (e.g., from IMU), as demonstrated by the ATE RMSE result of $6.91m$ reported in [peterson2025roman] on the same $hybrid$ data sequences. Furthermore, by experimental design, much of the trajectory overlap occured when robots were facing similar directions where visual feature matching methods excel.
@@ 7 Conclusions
While the results are promising, this work opens several directions for future research. For example, the robots are currently constrained to operating within the previously fused map, which limits the system from handling exploration tasks or adapting to changes in the environment. Further, the language grounding does not yet exploit the open-set nature of the object-based map to support more open-ended tasks. These are exciting directions for future research, with the potential to enable robots to complete a wider variety of tasks.